1. Let x and y represent two animal populations which satisfy (9-1-3y) y(-6 + 2x) (a) (5 points) What is the relationship bet(d) (7 points) Sketch the general solution. Be detailed. (e) (4 points) Sketch o(t) and y(t) if (0) = 1, y(0) = 5.
Answer #1

o civen system is. da dt - re (9-1-by) 위분위 dy at = y (-6 tare) a In the above system term ny denotes the relationship betweenrelationship is prey - predator relationship. nullclines. r(9-x-3y) y (-6+27) direction of systen. (0,3) (3,2) 2 |(3,0) remaiy will die out after some time. interior equilibrium in solution tollowing equations. g-x-3y =o -6+an - O → 2-3 putting 3 in4 General solution. y 4 3 1 2 co move in region 1, dk >o, dy dt at so in this region trajectories downwand and right in regio(3,2) stable critical point y (1,5) 16 (3,2)